Table salt and sugar are accidently mixed 5.50g sample is burned and a 2.90g of co2 is produced what is the mass percentage
1
Expert's answer
2014-11-18T12:23:10-0500
Only sugar takes partin this burning reaction. The molecular formula of sugar is C12H22011, so, the molar mass of sugar is 342g/mole.
The equation of this reaction is:
C12H22011+6O2=6CO2+11H2O
Suppose that x is number of grams of C12H22O11.
The molar mass of 6CO2is 264g/mole.
So, we get x/342 =2.90/264, x= (2.90*342)/264 = 3.76g.
The mass percentage of C12H22011 is (3.76/5.50)*100%=68.36%
The mass percentage of salt is 100% - 68.36% =31.64%